The width of a rectangle is 1 ft less than the length. The area is 2ft squared. Find the width and lenght.
Answer by RAY100(1637)

Area =Length * Width = 2 ft^2
.
but w = L-1
.
L * ( L-1) = 2
.
L^2 -L -2 = 0
.
( L -2) (L+1) =0
.
L-2 =0,,,,L=2,,,,and w=L-1 = 1,,,ok
.
or L+1 =0,,,,,L=-1,,,,which is not reasonable
.
check
.
Area = L*W = 2*1 =2 ,,,,ok